OpenStudy (anonymous):

Lines IJ and KL I(1,0) J(5,3) K(6,-1) L(0,2) Are the lines Perpendicular or Parallel?

OpenStudy (anonymous):

Slope = difference in y/difference in x Slope of IJ: (0-3)/(1-5) = -3/-4 = 3/4 Slope of KL: (-1-2)/(6-0) = -3/6 = -1/2 Neither; slopes would have to be equal to be parallel, or be the negative reciprocals to be perpendicular (ie, 2/3 and -3/2).